# BiFuzz: A Two-Stage Fuzzing Tool for Open-World Video Games

### v1

@@ Abstract
Open-world video games present a broader search space than other games, posing challenges for test automation. Fuzzing, which generates new inputs by mutating an initial input, is commonly used to uncover failures. In this study, we proposed BiFuzz, a two-stage fuzzer designed for automated testing of open-world video games, and investigated its effectiveness. The results revealed that BiFuzz mutated the overall strategy of gameplay and test cases, including actual movement paths, step by step. Consequently, BiFuzz can detect ‘stucking’ failures. The tool and its video are at https://github.com/Yusaku-Kato/BiFuzz.
Index Terms:
open-world video game, fuzzing
@@ I Introduction
Increasing costs and prolonged development periods are significant problems in video game development[1][2]. Automation is one of the ways to address prolonged development periods[3]. In video game development, several studies have been performed on automated testing[4][5]. Several game companies employ scripts to simulate player actions for test automation[6]. However, many developers remain reluctant to adopt test automation, as scripted tests are vulnerable to frequent software changes[6]. Video games undergo frequent modifications throughout development, often driven by evolving subjective and artistic requirements from game designers. These changes aim to ensure the game remains enjoyable for players[6]. Design-driven modifications can render automated tests obsolete, requiring continual updates to test content. Additionally, the cost of employing human testers for playtesting is currently lower than that of implementing test automation[6]. For video game test automation, the above problems need to be solved.
Automating testing is challenging, especially for open-world video games, allowing players to explore vast environments freely. In the games, players enjoy a great deal of autonomy in selecting from a complex system of mechanics, explorable spaces, and goals. The games feature a diverse array of objects (e.g., buildings and creatures) within a single scene. Furthermore, flexible movement mechanics and diverse object interactions enhance player freedom, contributing to an enormous state space that is challenging to explore[7]. The existing studies on automated playing in open-world video games have proposed LLM-based approaches[8][9]. However, these approaches require large volumes of training data. As video games undergo frequent modifications throughout development, training data must be collected after each change. Thus, employing LLM in game development is cost-prohibitive and impractical.
Fuzzing is a testing approach that repeatedly executes a program with automatically generated inputs—often syntactically or semantically unexpected—to detect failures[10]. Fuzzing requires no training data and adapts easily to frequent software changes. Accordingly, this study focuses on fuzzing for automated testing.
In this study, we propose BiFuzz, a two-stage fuzzing tool that considers state transitions in open-world video games. A summary of our contributions is as follows:
@@
  • We demonstrate that BiFuzz can successfully detect pre-inserted ‘stucking’ failures when applied to an open-world video game developed using Unity.
@@ II Related Work
Fuzzing is an automated testing approach, in which an initial input is repeatedly mutated to generate new inputs, called fuzz, that are likely to cause failures. In this study, a ‘character’ refers to an object controlled by the player in video games. ‘Fuzz’ denotes an input sequence used to move a character that is likely to cause failures. AFL [^1] https://lcamtuf.coredump.cx/afl/ is a representative coverage-guided grey-box fuzzer with a proven track record of detecting failures in various software. It measures code coverage for each input and uses this information to guide the generation of subsequent fuzz. Aschermann et al.[11] proposed IJON to enhance AFL and applied it to Super Mario Bros. They demonstrated that conventional fuzzing based on code coverage performs inadequately for software with complex state transitions, such as video games. Furthermore, they proposed an extended tool named IJON to improve AFL based on annotations. Consequently, they demonstrated the alleviation of state explosion.
Modern video games released are typically developed using game engines [12]. Unity, a widely used game engine, provides functions, such as object placement, physics calculations, and graphics drawing. Open-world video games are predominantly developed using such engines. Classical video games (e.g., Super Mario Bros.) have a few status variables. However, applying IJON to open-world video games developed using game engines is challenging due to the complexity of various status variables and the need to generate simultaneous inputs for movement and actions. BiFuzz is proposed to address these limitations in prior work.
Figure 2: Flowchart for determining $R_{t}$
@@ III Two-Stage Fuzzing Tool: BiFuzz
In open-world video games, a player has various in-game destinations, or keypoints and can freely change the routes for the keypoints. This flexibility contributes to a state explosion. To explore this state space efficiently, we developed a tool BiFuzz, which has a two-stage fuzzing process. According to SteamDB[12], Unity is the most widely used game engine. Therefore, we developed BiFuzz to operate on the Unity platform. BiFuzz was developed as a component within Unity, allowing its scripts to be attached to character object via drag-and-drop. This approach enables fuzzing for video games developed using game engines incompatible with IJON[11] and supports common open-world video games that require simultaneous inputs.
Fig. 1 provides an overview of fuzzing process in BiFuzz. The tool operates within the IDE (i.e., Unity Editor) and iteratively generates and mutates an initial input. BiFuzz performs fuzzing in two stages: a global fuzzing and a local fuzzing. Fig. 1 includes both stages of fuzzing. In the process of a global fuzzing, BiFuzz determines the order in which players pass through keypoints, which are represented as yellow stars. During a local fuzzing, waypoints—represented as red stars—are added between each pair of yellow stars determined in a global fuzzing. BiFuzz simulates player character movements by navigating through these waypoints, enabling detailed map exploration. BiFuzz repeats these two fuzzing stages until it processes the number of frames defined by a tester for the target video game. Let $m$ denote the number of iterations during the tester-defined time, and let $t$ be a natural number such that $t\leq m$. Let $R_{1},R_{2},...,R_{m}$ be the iteration results obtained from the 111st to the $m$-th iterations. Subsequently, each result $R_{t}$ and the corresponding detected failure $F_{t}$ are represented in the following format:
$F_{t}=(R_{t},C_{t}),\;R_{t}=Pass/Fail/Timeout$
$C_{t}$ is defined only when $R_{t}$ is $Fail$, indicating the coordinate of an occurred failure. $R_{t}$ is $Fail$ if BiFuzz detects a failure, and $Pass$ if the entire input sequence is executed without detecting any failure. Furthermore, to prevent excessive delays caused by a large input sequence, the value of $R_{t}$ is set to $Timeout$ if the elapsed time of an iteration exceeds a tester-defined threshold. Fig. 2 shows a flowchart of the determination of $R_{t}$ in BiFuzz. The termination condition for BiFuzz is a target video game reaching the number of frames specified by a tester. Furthermore, for each iteration, two cases are considered. BiFuzz selects the appropriate case based on the observation of $F_{t}$. The following outlines the two cases. If the result is the same as the previous iteration, BiFuzz proceeds to Case 1, executing both a global and a local fuzzing. Otherwise, it proceeds to Case 2, executing only a local fuzzing. The details of a global fuzzing and a local fuzzing are described below.
@@ III-A Global Fuzzing
A character is an object controlled by the player. BiFuzz imitates typical player operations by determining the order in which the character passes through keypoints. This order is guided by a parameter that records each keypoint’s pass/not pass status and priority. In this paper, a play style is defined by the characteristics of the player’s operation expected by testers, and is configured prior to fuzzing. BiFuzz assigns weighted probabilities to candidate parameter values based on the specified play style. For example, in a play style where the character is expected to pass through all keypoints, all boolean parameters indicating keypoint passage are set to $true$. As illustrated, parameter values are selected based on a weighted probability distribution.
@@
In this study, a strategy refers to the keypoints a character passes through and the order in which they are traversed. This paragraph describes the process of determining the strategy. Let keypoints be stored in an array $k=[1,2,...,n]$, where $n$ is the number of keypoints. Let $x$ be a natural number such that $x\leq n$, Keypoint $x$ exists at coordinate $l_{x}$. BiFuzz employs strategy-determining parameters in this process. Let parameters for each keypoint be denoted by $PG_{x}=(A_{x},B_{x})$. Here, $A_{x}$ is a boolean value that records the pass/not-pass status of Keypoint $x$, and $B_{x}$ is an integer between 0 and $n$ that represents its priority. As an example, consider a case where there are Keypoints 111, 222, and 333 in the game, and the parameters for each keypoint are $PG_{1}=(\top,0),PG_{2}=(\bot,0)$, and $PG_{3}=(\top,1)$. Fig. 3 shows the process of strategy determination under the above parameter settings. In this figure, as the parameter $A$ is $true$ for both $PG_{1}$ and $PG_{3}$ and $false$ for $PG_{2}$, the character aims for Keypoints 111 and 333. Furthermore, the parameter $B$, which indicates the priority level of Keypoint 333, is 111. Therefore, it is higher than Keypoint 111, which is 0. Consequently, the character aims for Keypoint 333 before Keypoint 111. Thus, the strategy in this figure becomes $Strategy=[l_{3},l_{1}]$.
@@ III-B Local Fuzzing
To add waypoints into the order determined in a global fuzzing, BiFuzz employs parameters in a consistent manner. Each parameter in a local fuzzing includes the number and coordinates of waypoints, and it is selected based on the probabilities weighted by the specified play style.
@@
In this study, BiFuzz determines a degree of freedom by adding waypoints to a strategy. This paragraph explains the process of incorporating waypoints into the strategy determined in a global fuzzing using waypoint-adding parameters. Let parameters for each keypoint be denoted by $PL_{x}=(C_{x},D_{x},E_{x})$. Here, $C_{x}$ is an integer between 0 and 999999 that represents the number of waypoints to be added between Keypoint $x$ and Keypoint $x+1$. $D_{x}$ is an integer between 0 and 999999 that represents the maximum distance from Keypoint $x$ to a waypoint. Moreover, $E_{x}$ is an integer between 111 and 444 that determines the direction of the waypoint from Keypoint $x$. In the same example used in the explanation of a global fuzzing, a new element is added to the array $Strategy=[l_{3},l_{1}]$. As no keypoint exists after Keypoint 111 in $Strategy$, only the parameters for $PL_{3}$ are used. Let $PL_{3}$ be defined as $PL_{3}=(1,50,1)$. Fig. 4 shows a top-down view of the map in Star Collection, an open-world video game developed using Unity. When $PL_{3}$ is set as above, an added waypoint is randomly selected from the light blue region shown in Fig. 4.
[…]
@@ IV Evaluation
We confirmed that BiFuzz performs effective testing. Specifically, we prepared two play styles for BiFuzz’s fuzzing, conducted fuzzing with each, and confirmed how much the number of failures detected could be increased. Star Collection is an open-world video game developed by the first author using Unity. We applied BiFuzz to one scenario in Star Collection. The scenario was made to confirm the effectiveness of BiFuzz for a ‘stucking’ failure commonly occurring in an open-world video game. In Star Collection, a player operates a character in third-person view, with the objective of collecting items located at keypoints. Character movement is controlled using the W, A, S, and D keys, while the viewpoint is adjusted with the mouse.
We present the results of the test for the scenario that includes ‘stucking’ failures. At first, we created a hundred objects for implementing ‘stucking’ failures, and indexed them from 1 to 100. Subsequently, we tested each of the Play styles A and B. In Play style A, a character passes through fewer keypoints and waypoints than Play style B. We applied BiFuzz to Star Collection, and executed it on the IDE (i.e., Unity Editor) for 2,592k frames (12 hours) with the target frame rate of 60 fps. We conducted it five times.
Figure 5: Number of detected failures
On average, 8.4 more scenes were executed during testing with Play style A than with Play style B. This is because Play style A involves fewer keypoints and waypoints, resulting in a relatively smaller input in one scene. Additionally, the number of detected failures indicates how many of the 100 predefined failures in the scenario were identified. In Play style A, an average of 2.8 more failures were detected than in Play style B. Fig. 5 presents a graph comparing the trends in the number of failures detected in Play styles A and B, with the average number of detected failures on the vertical axis and the number of executed scenes on the horizontal axis. The graph shows that Play style A detected more failures than Play style B, primarily due to the higher number of scenes executed during testing.
A total of 64 failures were detected using Play style A and 56 using Play style B. We analyzed the overlap and uniqueness of these failures across five runs. We found that, 43 failures were common to both styles, 21 were unique to Play style A, and 13 were unique to Play style B. Using different play styles enables the detection of distinct failures that may be missed by others.
@@ V Conclusion
We proposed BiFuzz, a two-stage fuzzing tool for testing open-world video games. The experimental results show that each of the two play styles detected unique failures. This indicates that BiFuzz can detect failures, and that performing fuzzing with two distinct play styles uncovers a greater number of failures.

### v2

@@ Abstract
Open-world video games present a broader search space than other video games, posing challenges for test automation. Fuzzing, which generates new inputs by mutating an initial input, is commonly used to uncover issues. In this study, we proposed BiFuzz, a two-stage fuzzer designed for automated testing of open-world video games, and investigated its effectiveness. The results revealed that BiFuzz mutated the overall strategy of gameplay and test cases, including actual movement paths, step by step. Consequently, BiFuzz can detect character stuck issues. The tool and its video are at https://github.com/Yusaku-Kato/BiFuzz and https://www.youtube.com/watch?v=VOrHfnLJSbk.
@@ I Introduction
Increasing costs and prolonged development periods are significant problems in video game development[Untitled27:online][2023shit96:online]. Automation is a way to address prolonged development periods[5387726]. In video game development, several studies have been performed on automated testing[prasetya2020navigation][10.1145/3550355.3552440]. Several video game companies employ scripts to simulate player actions to test automation[murphy2014cowboys]. However, many developers remain reluctant to adopt test automation, as scripted tests are vulnerable to frequent software changes[murphy2014cowboys]. Video games undergo frequent modifications throughout development, often driven by the game designers’ evolving subjective and artistic requirements. These changes aim to ensure the game remains enjoyable[murphy2014cowboys]. Design-driven modifications can render automated tests obsolete, requiring continual updates to test content. Additionally, the cost of employing human testers for playtesting is currently lower than that of implementing test automation[murphy2014cowboys]. For video game test automation, the above problems need to be solved.
It is challenging to automate testing, especially for open-world video games, which allow players to explore vast environments freely. In these games, players enjoy considerable autonomy in selecting from a complex system of mechanics, explorable spaces, and goals. The games feature a diverse array of objects (e.g., buildings and creatures) within a single scene. Furthermore, flexible movement mechanics and diverse object interactions enhance player freedom, contributing to an enormous state space that is challenging to explore[10.1145/3102071.3102098]. The existing studies on automated playing in open-world video games have proposed LLM-based approaches[wang2023voyager]. However, these approaches require large volumes of training data. As video games undergo frequent modifications throughout development, training data must be collected after each change. Thus, employing LLMs in video game development is cost-prohibitive and impractical.
To detect potential issues, the fuzzing testing approach repeatedly executes a program with automatically generated inputs that are often syntactically or semantically unexpected [8863940]. Fuzzing requires no training data and adapts easily to frequent software changes. Furthermore, fuzzing can reduce development effort by eliminating the processes in which testers write scripts for each test case or operate the game. Accordingly, this study focuses on fuzzing for automated testing.
This paper presents BiFuzz, a two-stage fuzzing tool designed specifically to detect character stuck issues (CSI), which are common in open-world video games. Our contributions are that:
@@
  • We demonstrate that BiFuzz can successfully detect pre-inserted CSIs when applied to an open-world video game developed using Unity.
@@ II Related Work
Fuzzing is an automated testing approach in which an initial input is repeatedly mutated to generate new inputs, termed fuzz, that are likely to cause issues. In this study, a ‘character’ refers to an object controlled by the player in video games. ‘Fuzz’ denotes an input sequence used to move a character that is likely to cause issues. AFL [^1] https://lcamtuf.coredump.cx/afl/ is a representative coverage-guided grey-box fuzzer with a proven track record of detecting issues in various software. It measures code coverage for each input and uses this information to guide the generation of subsequent fuzz. Aschermann et al.[aschermann2020IJON] proposed IJON to enhance AFL and applied it to Super Mario Bros. They demonstrated that conventional fuzzing based on code coverage performs inadequately for software with complex state transitions, such as video games. Furthermore, they proposed an extended tool named IJON to improve AFL based on annotations. Consequently, they demonstrated the alleviation of state explosion.
Modern video games released are typically developed using game engines [SteamDB]. Game engines such as Unity provide functions, e.g., object placement, physics calculations, and graphics drawing. Open-world video games are predominantly developed using such engines. Classical video games (e.g., Super Mario Bros.) have a small number of state variables, making them easy to identify and annotate for IJON to work. Conversely, open-world video games developed using game engines have a larger number of state variables than those in classical video games, and these variables are difficult for testers to identify. This problem is common across game engines. Therefore, it is challenging to apply IJON to the open-world video games.
@@ III Two-Stage Fuzzing Tool: BiFuzz
We developed BiFuzz, which uses keypoints on the map to narrow the search space without annotations. In open-world video games, a player has various in-game destinations, or keypoints, and can freely change the routes for the keypoints. This flexibility contributes to a state explosion. To explore this state space efficiently, BiFuzz implements a two-stage fuzzing process. According to SteamDB[SteamDB], Unity is the most widely used game engine. Therefore, we developed BiFuzz to operate on the Unity platform. We developed it as a component within Unity, allowing its scripts to be attached to a character object via drag-and-drop. This approach enables fuzzing for open-world video games developed using game engines incompatible with IJON[aschermann2020IJON].
Fig. 1 provides an overview of the BiFuzz fuzzing process. The tool operates within the IDE (i.e., Unity Editor) and iteratively generates and mutates an initial input. BiFuzz performs fuzzing in two stages: Global Fuzzing (GF) and Local Fuzzing (LF). Fig. 1 includes both stages of fuzzing. In the GF process, BiFuzz determines the order in which players pass through keypoints, which are represented as yellow stars. During LF, waypoints, represented as red stars, are added between each pair of yellow stars determined in GF. BiFuzz simulates player character movements by navigating through these waypoints, enabling detailed map exploration. In addition, execution traces and outcomes are recorded so that the generated inputs can be reproduced. BiFuzz repeats these two fuzzing stages until it processes the number of frames defined by a tester for the target video game. Let $m$ denote the number of iterations during the tester-defined time, and let $t$ be a natural number such that $t\leq m$. Let $R_{1},R_{2},...,R_{m}$ be the iteration results obtained from the 11st to the $m$-th iterations. Subsequently, each result $R_{t}$ and the corresponding detected issues $F_{t}$ are represented as $F_{t}=(R_{t},C_{t})$, where $R_{t}=Pass/Fail/Timeout$. $C_{t}$ is defined only when $R_{t}$ is $Fail$, indicating the coordinate of an occurred issue. $R_{t}$ is $Fail$ if BiFuzz detects an issue, and $Pass$ if the entire input sequence is executed without detecting any issues. Furthermore, to prevent excessive delays caused by a large input sequence, the value of $R_{t}$ is set to $Timeout$ if the elapsed time of an iteration exceeds a tester-defined threshold. The termination condition for BiFuzz is when a target video game reaches the number of frames specified by a tester. Furthermore, each iteration considers two cases. BiFuzz selects the appropriate case based on observing the $F_{t}$. The following outlines the two cases. If the result matches that of the previous iteration, BiFuzz proceeds to Case 1, executing both GF and LF. Otherwise, it proceeds to Case 2, executing only LF. The details of GF and LF are described below.
@@ III-A Global Fuzzing
A character is an object controlled by the player. BiFuzz imitates typical player operations by determining the order in which the character passes through keypoints. This order is guided by a parameter that records each keypoint’s pass/not pass status and priority. This paper defines a play style by the player operation characteristics expected by testers, and is configured prior to fuzzing. BiFuzz assigns weighted probabilities to candidate parameter values based on the specified play style. For example, in a play style where the character is expected to pass through all keypoints, all boolean parameters indicating keypoint passage are set to truetrue. As illustrated, the parameter values are selected based on a weighted probability distribution.
@@
In this study, a strategy refers to the keypoints a character passes through and the order in which they are traversed. This paragraph describes the process of determining the strategy. Let keypoints be stored in an array k=[1,2,,n]k=[1,2,...,n], where $n$ is the number of keypoints. Let $x$ be a natural number such that $x\leq n$, Keypoint $x$ exists at coordinate $l_{x}$. BiFuzz employs strategy-determining parameters in this process. Let the parameters for each keypoint be denoted by $PG_{x}=(A_{x},B_{x})$. Here, $A_{x}$ is a boolean value that records the pass/not-pass status of Keypoint $x$, and $B_{x}$ is an integer between 0 and $n$ that represents its priority. As an example, consider the case with Keypoints $1$, 22, and $3$ in a game, and the parameters for each keypoint are $PG_{1}=(\top,0),PG_{2}=(\bot,0)$, and $PG_{3}=(\top,1)$. Fig. 2 shows the process of strategy determination under the above parameter settings. In this figure, as the parameter $A$ is truetrue for both $PG_{1}$ and $PG_{3}$ and falsefalse for $PG_{2}$, the character aims for Keypoints $1$ and $3$. Furthermore, the parameter $B$, which indicates the priority level of Keypoint $3$, is $1$. Therefore, it is higher than Keypoint $1$, which is 0. Consequently, the character aims for Keypoint $3$ before Keypoint $1$. Thus, the strategy in this figure becomes $Strategy=[l_{3},l_{1}]$.
@@ III-B Local Fuzzing
To add waypoints into the order determined in GF, BiFuzz employs parameters in a consistent manner. Each parameter in LF includes the number and coordinates of waypoints, and it is selected based on the probabilities weighted by the specified play style.
@@
In this study, BiFuzz determines a degree of freedom by adding waypoints to a strategy. This paragraph explains the process of incorporating waypoints into the strategy determined in GF using waypoint-adding parameters. Let parameters for each keypoint be denoted by $PL_{x}=(C_{x},D_{x},E_{x})$. Here, $C_{x}$ is an integer between 0 and $99$ that represents the number of waypoints to be added between Keypoint $x$ and the next Keypoint that a character passes through. $D_{x}$ is an integer between 0 and $99$ that represents the maximum distance from Keypoint $x$ to a waypoint. Moreover, $E_{x}$ is an integer between $1$ and $4$ that determines the direction of the waypoint from Keypoint $x$. In the same GF explanatory example, a new element is added to the array $Strategy=[l_{3},l_{1}]$. As no keypoint exists after Keypoint $1$ in $Strategy$, only the parameters for $PL_{3}$ are used. Let $PL_{3}$ be defined as $PL_{3}=(1,50,1)$. Fig. 3 shows a top-down view of the map in Star Collection, an open-world video game developed using Unity. When $PL_{3}$ is set as above, an added waypoint is randomly selected from the light blue region shown in Fig. 3.
Next, based on Fig. 3, we explain how the degrees of freedom are determined using parameters. First, as the value of parameter $C_{3}$ is $1$, BiFuzz adds one waypoint. Moreover, from $E_{3}=1$, the added waypoint lies forward-right, assuming the positive Z-axis represents the front from the character’s perspective. In addition, as $D_{3}=50$, the maximum distance from Keypoint $3$ to the waypoint is set to half ($50\%$) of the minimum distance from Keypoint $3$ to each side of the map. This approach enables BiFuzz to explore the state space more effectively.
@@ IV Evaluation
Open-world video games generally feature distinct input types, such as simultaneous inputs and inputs of arbitrary length with arbitrary intervals. It is difficult to compare IJON, the existing fuzzer for video games, with BiFuzz because it does not support these types of inputs. To this end, we evaluated the following experiments to confirm that BiFuzz performs effective testing. Specifically, we prepared two play styles for the fuzzing, conducted fuzzing with each, and confirmed how much the number of issues detected could be increased. Star Collection is an open-world video game developed by the first author using Unity. We applied BiFuzz to one scenario in the game. The scenario was developed to confirm the effectiveness of BiFuzz for addressing the CSIs that commonly occur in open-world video games. The issue occurs when its movement is blocked due to colliding with the terrain or surrounding objects, hindering progress in the game. In the game, a player operates a character in third-person view, with the objective of collecting items located at keypoints. Character movement is controlled using the W, A, S, and D keys, while the viewpoint is adjusted with the mouse.
We present the results of the test for the scenario that includes CSIs. First, we created a hundred objects for implementing the issues, and indexed them from 1 to 100. Then, we tested each of the Play styles A and B. In Play style A (PS-A), a character passes through fewer keypoints and waypoints than in Play style B (PS-B). We applied BiFuzz to the game, and executed it on the IDE (i.e., Unity Editor) for 2,592k frames (12 hrs.) with the target frame rate of 60 fps. We conducted this routine five times.
Figure 4: Number of detected issues
On average, 8.4 more scenes were executed during testing with PS-A than with PS-B. This is because PS-A involves fewer keypoints and waypoints, resulting in a relatively smaller input in one scene. Additionally, the number of detected issues indicates how many of the 100 predefined issues in the scenario were identified. In PS-A, an average of 2.8 more issues were detected than in PS-B. Fig. 4 presents a graph comparing the trends in the number of issues detected in PS-A and PS-B, with the average number detected on the vertical axis and the number of executed scenes on the horizontal axis. The graph shows that PS-A detected more than PS-B, primarily due to the higher number of scenes executed during testing. A total of 64 issues were detected using PS-A and 56 using PS-B. We analyzed the overlap and uniqueness of these issues across five runs. We found that, 43 issues were common to both styles, 21 were unique to PS-A, and 13 were unique to PS-B. Using different play styles enables the detection of distinct issues that may be missed by others.
@@ V Conclusion
We propose BiFuzz, a two-stage fuzzing tool designed specifically to detect CSIs. The experimental results show that each of the two play styles detected unique issues. This indicates that BiFuzz can detect issues, and that performing fuzzing with two distinct play styles uncovers a greater number of CSIs.
However, the evaluation of BiFuzz was limited to a single custom game scenario, Star Collection, which makes the scale of the experiment insufficient. Therefore, in future work, we plan to establish baselines drawing on existing methods and conduct comparative evaluations with BiFuzz. Moreover, BiFuzz currently supports movement on the planar axes using the W, A, S, and D keys, and is unsupported for other input methods or three-dimensional movement. Nevertheless, it is technically feasible to extend the implementation to accommodate these in the future. In addition, since the experiments in this paper focused only on CSIs, future evaluations should also include other common video game issues, including graphics rendering errors and performance issues, to enhance the generality of the conclusions.
Acknowledgment
This work was supported by JSPS KAKENHI Grant Number JP24K02923 and JST AIP Acceleration Research JPMJCR25U7, Japan.